Construct the discrete series. Count unbiased estimates of the general mean and general variance. Find the confidence interval for expectation with a confidence level of 0.05 5, 11, 13, 9, 11, 5, 7, 7, 5, 9, 13, 13, 11, 9, 5, 9, 11, 9, 5, 9, 11, 9, 9, 5.


1
Expert's answer
2021-12-30T06:26:01-0500

(i)

"n=24"


"\bar{x}=\dfrac{1}{24}(5+11+13+9+11+5+7+7"

"+5+9+13+13+11+9+5+9+11+9"

"+5+9+11+9+9+5)=8.75"

Unbiased estimates of the general mean is "8.75."



"Var(X)=s^2=\dfrac{1}{24-1}((5-8.75)^2+(11-8.75)^2"

"+(13-8.75)^2+(9-8.75)^2+(11-8.75)^2"

"+(5-8.75)^2+(7-8.75)^2+(7-8.75)^2"

"+(5-8.75)^2+(9-8.75)^2+(13-8.75)^2"

"+(13-8.75)^2+(11-8.75)^2+(9-8.75)^2"

"+(5-8.75)^2+(9-8.75)^2+(11-8.75)^2"

"+(9-8.75)^2+(5-8.75)^2+(9-8.75)^2"

"+(11-8.75)^2+(95-8.75)^2+(9-8.75)^2"

"+(5-8.75)^2)=7.413043"

Unbiased estimates of the general variance is "7.413043."

"s=\sqrt{s^2}=2.7227"

(ii)The critical value for "\alpha = 0.05" and "df = n-1 = 23"  degrees of freedom is "t_c = z_{1-\alpha/2; n-1} = 2.068658."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}},\bar{x}+t_c\times\dfrac{s}{\sqrt{n}})"

"=(8.75-2.068658\times\dfrac{2.7227}{\sqrt{24}},"

"8.75+2.068658\times\dfrac{2.7227}{\sqrt{24}})"

"=(7.6003, 9.8997)"

Therefore, based on the data provided, the 95% confidence interval for the population mean is "7.6003 < \mu < 9.8997," which indicates that we are 95% confident that the true population mean "\mu" is contained by the interval "(7.6003,9.8997)."